# Global Public Sentiment on Decentralized Finance: A Spatiotemporal Analysis of Geo-tagged Tweets from 150 Countries

### v1

@@ Abstract
In the digital era, blockchain technology, cryptocurrencies, and non-fungible tokens (NFTs) have revolutionized financial and decentralized systems. Despite their importance, current research frequently overlooks the spatiotemporal variations in public sentiment toward these technologies, thereby limiting the ability to draw macro-level inferences on how these technologies have influenced global public attention, governmental regulations, and economic growth. Recognizing the significance of these variations, this study utilizes Twitter as a primary data source to investigate public attention and sentiment across 150 countries. We analyze over 150 million geo-tagged tweets from the period 2012 to 2022, along with sentiment scores derived from a BERT-based multilingual sentiment classification model initially trained on 7.4 billion geotagged tweets. The study further integrates this data with global cryptocurrency regulation policies and national economic indicators from the World Development Indicators database. Our results uncover significant global variations in sentiment, influenced by economic factors such as GDP per capita, with more developed nations showing greater engagement in discussions, while less developed countries exhibit higher sentiment levels. A geographically weighted regression analysis confirms that the correlation between GDP and tweet engagement intensifies following Bitcoin price surges. Topic modeling reveals that countries within similar economic clusters share comparable discussion trends, whereas diverse clusters focus on different topics. This study underscores the global disparities in public sentiment towards decentralized finance, influenced by economic and regional factors. The implications of these findings are discussed concerning poverty alleviation, cryptocurrency-related crime, and sustainable development. We have made our dataset and code publicly available as open-source on GitHub.
@@
Blockchain, Cryptocurrency, Social Media, Twitter, Sentiment Analysis, Geographical Index
@@ I Introduction
In the digital age, the advent of blockchain technology, cryptocurrencies, and non-fungible tokens (NFTs) has fundamentally transformed our understanding of finance, trust, and decentralized systems. As these technologies continue to evolve and penetrate various sectors, understanding public sentiment towards them becomes increasingly important. Social media platforms, particularly Twitter, provide a valuable source of data for gauging public opinion and sentiment.
Previous research has primarily focused on how public sentiment can forecast the value of cryptocurrencies [1] and NFTs [2], as well as the differing attitudes toward decentralized finance among various groups, such as Key Opinion Leaders (KOLs) and general users [3].
However, there is a notable gap in research from a spatial analysis perspective. Few studies have explored how public attention and sentiment towards decentralized finance vary across different countries and regions. This exploration is crucial for uncovering global disparities in the understanding and engagement with these technologies. The absence of spatially oriented research represents a missed opportunity to grasp the nuanced ways in which different cultures and geopolitical landscapes shape public opinion and discourse on emerging digital financial technologies.
@@
The novelty of our research lies in the spatial analysis of geographically tagged Twitter data, which enables us to incorporate public attention and sentiment from both different timeframes and countries into our study. This approach allows us to identify how regional factors, cultural nuances, and economic conditions influence public attention and sentiment toward decentralized finance.
@@
In this study, we aim to investigate the following research questions (RQs):
@@
RQ1: How do public attention and sentiments expressed on social media toward decentralized finance vary across different countries and regions?
@@
RQ2: How do spatial variations in public attention and sentiment interact with economic factors, and how do these dynamics evolve over time?
RQ3: What specific topics do different countries and regions focus on when discussing decentralized finance?
@@
To address RQ1, we constructed a dictionary comprising 152 multilingual keywords to identify tweets related to decentralized finance, including blockchain, NFTs, and cryptocurrencies. Through data aggregation, we calculated sentiment scores and the proportion of decentralized finance-related tweets across approximately 150 countries from 2012 to 2022. Our preliminary findings reveal significant global variations in both the level of attention to decentralized finance and the sentiments expressed toward these topics.
@@
To address RQ2, we introduced GDP per capita as a proxy for the economic development of different countries. Through exploratory analysis, we examined the correlations among GDP per capita, tweet proportion, and sentiment score to investigate how economic factors influence public attention and sentiment towards decentralized finance across various countries. We then conducted a geographically weighted regression analysis between GDP per capita and the proportion of tweets concerning decentralized finance for each year. Our findings reveal that GDP per capita becomes a significant predictor of tweet proportion from 2015 onwards, highlighting the role of economic development in driving social media engagement with new technologies. More economically developed countries show higher levels of engagement, while less developed nations exhibit lower levels. To assess changes over time, we performed a Pearson correlation analysis between GDP per capita and tweet proportion for each year, followed by Fisher’s z transformation of the correlation coefficients. The results indicate that the influence of GDP on attention to decentralized finance fluctuates over time, often aligning with changes in the price of digital currencies, such as Bitcoin. Notably, after significant increases in Bitcoin’s value, the impact of GDP on attention levels tends to strengthen, underscoring the sensitivity of economically developed countries to shifts in economic value. Furthermore, countries with lower economic development but high engagement in decentralized finance often display the highest sentiment levels. Based on these three variables—GDP per capita, tweet proportion, and sentiment score—we conducted a cluster analysis, identifying three distinct clusters of countries.
To address RQ3, we conducted topic modeling on tweets from each of the three clusters of countries. After classifying 150 countries into three clusters based on key factors, we identified distinct differences in decentralized finance topics discussed among these clusters. Our analysis reveals that both lowest-GDP and mid-GDP countries emphasize the use of cryptocurrency to improve their economic situations. However, the lowest-GDP countries often view cryptocurrency as a means of achieving sudden wealth, as indicated by terms like ”luck,” ”hope,” and ”promote.” Mid-GDP countries, on the other hand, associate decentralized finance with social and religious elements, using terms such as ”people,” ”god,” and ”lord.” In contrast, the highest-GDP countries perceive cryptocurrencies as innovative financial tools for entertainment and daily life enhancement.
@@ V Results
Figure 1: Proportions of tweets related to decentralized finance across countries
Figure 2: Sentiment scores of tweets related to decentralized finance across countries
@@ V-A Spatial Autocorrelation
The results of Table I elucidate that our variables and clusters suffer spatial autocorrelation since all the p-values are relatively small. Therefore, we should utilize the geographically weighted regression instead of simple linear regression to conduct the regression analysis.
@@
Note: Standard errors are presented in parentheses. * p $<$ 0.05, ** p $<$ 0.01, *** p $<$ 0.001.
@@ V-B Geographically Weighted Regression
Note: Standard errors are presented in parentheses. * p $<$ 0.05, ** p $<$ 0.01, *** p $<$ 0.001.
GDP per capita represents the value for each country in each year. The tweet proportion is determined by the volume of tweets related to decentralized finance, normalized by the total volume of tweets in each country for each year. Both variables are standardized before regressions.
Figure 3: Fisher’s transformation of correlation coefficients across years
@@ V-C Fisher’s Z Transformation of Correlation
Figure 3 concludes the results of robustness checks. We dedicate to testing whether GDP per capita still affects the proportion of tweets related to specific keywords of decentralized finance instead of the entirety. It evinces that when we test the GDP per capita by introducing general keywords, such as “blockchain,” “NFT,” “Cryto,” the result maintains that it positively impacts the proportion of tweets related to general keywords after 2015. However, the GDP per capita is not statistically significant in affecting the proportion of tweets related to certain specific keywords, such as Ethereum.
@@ V-D Spatial Clustering
Figure 4: Global clustering of countries
Figure 5: Scatter plot of GDP per capita, sentiment score, and proportion of tweets related to decentralized finance
Figure 6: Box plot of GDP per capita, sentiment score, and proportion of tweets related to DeFi for each cluster of countries
@@
Cluster 1 Cluster 2 Cluster 3
GDP per capita (USD) Mean 3075.595 4222.720 33879.133
Median 2256.593 2748.896 26785.457
Min 240.087 399.657 4099.107
Max 11795.697 20064.097 117302.843
Std 2811.690 4203.290 25430.121
Tweet proportion (%) Mean 0.497 0.102 0.280
@@
Std 0.470 0.075 0.186
Sentiment score Mean 0.706 0.651 0.665
@@ V-D1 Descriptive Statistics
Countries in Cluster 1 and Cluster 2 are characterized by relatively underdeveloped economies, with average GDP per capita values of \$3,075.595 and \$4,222.720, respectively. Our regression results in Table II indicate that countries with more advanced economic development tend to have higher public attention towards decentralized finance. However, countries in Cluster 1 present an interesting anomaly, exhibiting the highest level of attention to decentralized finance on social media, as reflected by the average proportion of related tweets (0.497%) and the most positive sentiment (average sentiment score of 0.706). In contrast, countries in Cluster 2 demonstrate the lowest level of attention to decentralized finance (average proportion of related tweets of 0.102%) and a relatively low sentiment (average sentiment score of 0.651).
Countries in Cluster 3, which have the most developed economies (average GDP per capita of \$33,879.133), show levels of attention to decentralized finance (average proportion of related tweets of 0.280%) and sentiment scores (average sentiment score of 0.665) that are lower than those in Cluster 1 but higher than those in Cluster 2. Figure 6 presents box plots depicting various countries based on the three variables mentioned above.
@@ V-D2 Crypto Legal Status
Considering the varying legal statuses of cryptocurrency trading in different countries, which may influence the level of attention and sentiment towards topics related to decentralized finance, we incorporated the “Crypto Legal Status” data from the “Cryptocurrency Regulation Tracker”. Box plots (Figure 7A) reveal a correlation between GDP per capita and the legality of cryptocurrencies. Excluding the outliers, South Korea and China, countries that recognize the legality of cryptocurrencies, those with a partial ban, and those with a general ban on cryptocurrencies show decreasing levels of economic development, with average GDP per capita values of \$24,009.788, \$7,635.754, and \$3,565.186, respectively. We assigned values to the three crypto legal statuses (Legal, Partial Ban, General Ban) as 0, 0.5, and 1, respectively, to represent the degree of prohibition of cryptocurrencies. The Spearman correlation coefficient between this variable and GDP per capita is -0.530 (p-value $<$ 0.001). The proportions of countries within different clusters that recognize the legality of cryptocurrency (Figure 7B) also show a similar pattern. In the most economically developed Cluster 3, 30% (15 out of 50) of the countries recognize the legality of cryptocurrency trading. In contrast, only 15% (5 out of 33) in Cluster 1 and 11% (7 out of 62) in Cluster 2 do so.
Figure 7: Crypto Legal Status
V-E Topic Modelling
Figure 8: Topic probability for each cluster across years
Figure 9: Topic probability for each cluster across years
@@ V-E Topic Modelling
Both lowest-GDP and mid-GDP countries emphasize the use of cryptocurrency to improve their economic situations, but the lowest-GDP countries surprisingly may consider cryptocurrency as a tool to achieve sudden wealth, characterized by words like ”luck,” ”hope,” and ”promote.” The mid-GDP countries focus on connecting social and religious elements with decentralized finance, using terms such as ”people,” ”god,” and ”lord.” However, the highest-GDP countries regard cryptocurrencies as novel financial tools to entertain and supplement daily life.

### v2

@@ Abstract
Blockchain technology and decentralized finance (DeFi) are reshaping global financial systems, yet existing research often overlooks the spatial distribution of public sentiment and its economic and geopolitical determinants. To address this gap, we analyzed over 150 million geo-tagged, DeFi-related tweets from 2012 to 2022, drawn from a comprehensive dataset of 7.4 billion tweets. Employing sentiment scores generated by a BERT-based multilingual classification model, we integrated these tweets with economic and geopolitical data to construct a multimodal dataset. We utilized techniques such as sentiment analysis, spatial econometrics, clustering, and topic modeling to elucidate the drivers of engagement and sentiment. Our results reveal significant global variations in attention to DeFi and the sentiments expressed, with economic development emerging as a pivotal influence, particularly after 2015. Geographically weighted regression analysis confirms GDP per capita as a key predictor of the proportion of DeFi-related tweets, with its influence intensifying following notable increases in cryptocurrency values, such as Bitcoin. Wealthier nations generally exhibit more active engagement in DeFi discourse, while the lowest-income countries, displaying high attention and positive sentiments, frequently frame DeFi in terms of financial security and potential for sudden wealth. In contrast, middle-income countries contextualize DeFi within social and religious narratives, whereas high-income countries predominantly regard it as a speculative instrument or a form of entertainment. This study not only advances interdisciplinary research in computational social science and finance but also contributes to open science by releasing our dataset and analytical code on GitHub, and by providing a non-code workflow on the KNIME platform. These resources empower a diverse range of scholars to investigate DeFi adoption and sentiment, thereby informing policymakers, regulators, and developers in promoting financial inclusion and fostering responsible DeFi engagement worldwide.
@@
Blockchain, Decentralized Finance, Cryptocurrency, Social Media, Twitter, Sentiment Analysis, Geo-spatial Economics
@@ I Introduction
The emergence of blockchain technology, cryptocurrencies, and non-fungible tokens (NFTs) has fundamentally reshaped our understanding of finance, trust, and decentralized systems [1, 2]. Among these innovations, decentralized finance (DeFi) has become a disruptive force, challenging traditional financial intermediaries by providing open, permissionless, and algorithm-driven financial services [3, 4]. As DeFi continues to expand and integrate into the global economy, it is essential to understand the factors driving its adoption and impact.
One critical determinant of DeFi adoption is public sentiment, which significantly influences how emerging financial technologies are embraced and incorporated into the broader economy [5, 6]. Research on the adoption of other emerging technologies, such as artificial intelligence (AI), has demonstrated that positive sentiment fosters higher adoption rates and investment flows, whereas negative perceptions contribute to hesitancy, regulatory scrutiny, and reduced participation [7, 8]. Social media platforms, particularly Twitter, serve as valuable data sources for tracking public opinion and sentiment, providing real-time insights into market trends and user perceptions [9]. Previous studies on sentiment toward decentralized finance have primarily examined its role in forecasting the value of cryptocurrencies [10] and NFTs [11], as well as the differing attitudes toward DeFi among various groups, such as Key Opinion Leaders (KOLs) and general users [12]. However, there remains a notable gap in research from a spatial analysis perspective. Few studies have investigated how public attention and sentiment toward DeFi vary across different countries and regions. Exploring these variations is crucial for understanding global disparities in the perception and engagement with decentralized financial technologies, which would, in turn, affect the adoption of the emerging technology DeFi and thus influence economic development. The lack of spatially oriented research represents a missed opportunity to assess how economic, regional, and geopolitical factors shape public discourse and sentiment regarding emerging digital financial systems.
@@
Our research is unique in its spatial analysis of geographically tagged Twitter data, which enables the integration of public attention and sentiment across different timeframes and countries. This approach allows us to identify how regional economic conditions and geopolitical contexts influence public sentiment toward DeFi, offering new insights into the global adoption of decentralized financial technologies.
@@
As an initial exploratory step to pave the way for further causal hypothesis testing, this study aims to investigate the following research questions (RQs):
@@
RQ1: Are there statistically significant variations in public attention and sentiments expressed on social media toward decentralized finance across different countries and regions? If so, what are the key differences?
@@
RQ2: Are spatial variations in public attention and sentiment significantly associated with economic factors, and how do these dynamics evolve over time?
RQ3: What specific topics are countries and regions most likely to discuss when engaging with decentralized finance, and how do these topics differ across clusters formed based on economic, public attention, and sentiment differences?
@@
To address RQ1, we compiled a dictionary of 152 multilingual keywords to identify tweets related to decentralized finance, including blockchain, NFTs, and cryptocurrencies. By aggregating data, we computed sentiment scores and the proportion of decentralized finance-related tweets across approximately 150 countries from 2012 to 2022. Our preliminary findings indicate significant global variations in both the level of attention to decentralized finance—particularly when controlling for general engagement on Twitter—and the sentiments expressed toward these emerging topics.
@@
To address RQ2, we introduced GDP per capita as a proxy for the economic development of different countries. Through exploratory analysis, we examined the correlations among GDP per capita, tweet proportion, and sentiment score to investigate how economic factors influence public attention and sentiment towards decentralized finance across various countries. We then conducted a geographically weighted regression analysis between GDP per capita and the proportion of tweets concerning decentralized finance for each year. Our findings reveal that GDP per capita becomes a significant predictor of tweet proportion from 2015 onwards, highlighting the role of economic development in driving social media engagement with new technologies. Controlling for general engagement on Twitter, we find that more economically developed countries show higher levels of engagement on DeFi, while less developed nations exhibit lower levels. To assess changes over time, we performed a Pearson correlation analysis between GDP per capita and tweet proportion for each year, followed by Fisher’s z transformation of the correlation coefficients. The results indicate that the influence of GDP on attention to decentralized finance fluctuates over time, often aligning with changes in the price of digital currencies, such as Bitcoin. Notably, after significant increases in Bitcoin’s value, the impact of GDP on attention levels tends to strengthen, underscoring the sensitivity of economically developed countries to shifts in economic value. Furthermore, countries with lower economic development but high engagement in decentralized finance often display the highest sentiment levels.
To address RQ3, we applied clustering analysis, data visualization, and topic modeling to examine tweets from each of the three identified groups of countries. The clustering analysis categorized countries based on GDP per capita, public attention toward decentralized finance, and sentiment expressed on social media, yielding an optimal Silhouette Score of 0.312. While higher-income countries generally exhibited greater public attention to decentralized finance, lower-income countries presented an anomaly, demonstrating the highest levels of engagement and the most positive sentiment, whereas middle-income countries displayed the lowest levels of both attention and sentiment. Furthermore, with the exception of a few outliers such as China and South Korea, countries where cryptocurrency is legally recognized tend to have higher GDP per capita, and the legal status of cryptocurrencies follows a pattern in which economically advanced nations are more likely to permit cryptocurrency trading. The subsequent topic analysis provides insight into this anomaly, revealing that discussions in lower-income countries are more emotionally charged and frequently frame decentralized finance as a means of achieving financial security or sudden wealth, as indicated by terms such as “luck,” “hope,” and “promote.” In contrast, middle-income countries emphasize the intersection of decentralized finance with social and religious contexts, while high-income countries primarily regard cryptocurrencies as financial instruments for investment or entertainment.
@@ V Results
Figure 1: Proportions of tweets related to decentralized finance across countries. The figure illustrates the spatial distribution of tweets containing at least one decentralized finance-related keyword. Darker shades indicate a higher proportion of relevant tweets originating from that country, while lighter shades represent lower proportions.
Figure 2: Sentiment scores of tweets related to decentralized finance across countries. The figure visualizes the average sentiment score of tweets containing at least one decentralized finance-related keyword. Higher sentiment scores (in darker shades) indicate a more positive sentiment, while lower scores (in lighter shades) reflect a more negative sentiment.
@@ V-A Spatial Autocorrelation
Figures 1 and 2 provide a geographic overview of decentralized finance discussions on Twitter. The first map highlights the proportion of relevant tweets per country, while the second visualizes their sentiment distribution.
The results in Table I elucidate that our variables and clusters exhibit spatial autocorrelation, as indicated by the relatively small p-values. This suggests that decentralized finance-related discussions are not randomly distributed across space but tend to cluster geographically. Consequently, a geographically weighted regression (GWR) is more appropriate than a simple linear regression for analyzing these patterns, as GWR accounts for spatial heterogeneity and localized variations in relationships between variables.
@@
Note: This table presents the results of the spatial autocorrelation analysis using Moran’s I and Chi-square tests across different spatial weight structures. Moran’s I for GDP per capita, tweet proportion, and sentiment score: The positive and statistically significant Moran’s I values indicate strong spatial clustering, meaning that similar values of these variables tend to be geographically concentrated. The effect is strongest for GDP per capita, suggesting that wealthier countries tend to cluster together. The spatial autocorrelation for tweet proportion and sentiment score, while weaker, remains significant across different nearest-neighbor specifications. Chi-square tests for country clusters: The spatial distribution of country clusters is tested using Chi-square statistics. Higher values indicate stronger clustering patterns. Cluster 3 (high-GDP countries) exhibits the highest level of spatial clustering, followed by Cluster 2 and then Cluster 1. This suggests that economic and social factors influencing decentralized finance discussions are not randomly distributed but form distinct spatial patterns. Nearest-neighbor approach: The table reports results using 1 to 5 nearest neighbors to define spatial relationships. The decreasing magnitude of Moran’s I values as more distant neighbors are included suggests that spatial autocorrelation weakens with distance. Statistical significance levels: $p$ $<$ 0.05 (*), $p$ $<$ 0.01 (**), and $p$ $<$ 0.001 (***) indicate the robustness of these spatial effects. These findings emphasize that economic and social characteristics influencing decentralized finance discussions exhibit strong spatial dependencies. This reinforces the need for geographically weighted models when analyzing these trends.
@@ V-B Geographically Weighted Regression
Note: Standard errors are presented in parentheses. * $p$ $<$ 0.05, ** $p$ $<$ 0.01, *** $p$ $<$ 0.001. GDP per capita represents the economic output per person for each country in each year, measured in logarithmic form to adjust for skewness in income distribution. Tweet proportion represents the volume of tweets related to decentralized finance, normalized by the total volume of tweets in each country for each year. Spatial Lag ($\lambda$) measures the degree of spatial dependence in the regression model. A positive $\lambda$ suggests that decentralized finance tweet proportions in a given country are influenced by neighboring countries’ tweet proportions, indicating regional clustering effects. Pseudo $R^{2}$ is used to assess the model’s explanatory power. The increase in $R^{2}$ over the years suggests a growing relationship between GDP per capita and decentralized finance discussions.Key Observations: Before 2015, GDP per capita had an inconsistent effect on the proportion of decentralized finance-related tweets, with some years showing negative or insignificant relationships; Starting in 2015, the effect of GDP per capita became statistically significant, indicating that economic development increasingly influenced decentralized finance discussions; The highest coefficients occur between 2016 and 2018, suggesting a peak period where wealthier countries saw greater public engagement in decentralized finance discussions on Twitter; The spatial lag coefficient ($\lambda$) grows stronger in later years, suggesting that regional spillover effects in decentralized finance discussions became more pronounced. These findings highlight the evolving relationship between economic development and decentralized finance discussions and emphasize the role of spatial dependencies in shaping global Twitter activity trends.
Figure 3: Fisher’s transformation of correlation coefficients across years. The figure illustrates the temporal evolution of the correlation between GDP per capita and the proportion of tweets containing keywords related to decentralized finance. The Fisher Z-transformed correlation coefficients provide a stabilized measure of correlation, accounting for variations in sample size across years. Positive values indicate a stronger association between GDP per capita and keyword prevalence, while negative values suggest an inverse relationship. The confidence intervals represent the statistical uncertainty in estimating these correlations over time.
@@ V-C Fisher’s Z Transformation of Correlation
Figure 3 presents the results of our robustness checks, which examine the relationship between GDP per capita and the proportion of tweets related to decentralized finance keywords over time. Instead of analyzing the total volume of related tweets, we focus on specific keyword groups to assess whether economic factors influence discussions differently depending on the terminology used.
The results indicate that when considering general keywords such as blockchain, NFT, and crypto, GDP per capita maintains a positive and statistically significant correlation with the proportion of tweets containing these keywords, particularly after 2015. This suggests that higher-income countries tend to have a greater prevalence of general decentralized finance discussions on social media.
However, when narrowing the analysis to more specific terms such as Ethereum, GDP per capita does not exhibit a statistically significant effect on the proportion of tweets. This discrepancy suggests that while economic development may influence general awareness and engagement with decentralized finance, discussions around specific projects or technologies may be driven by other factors, such as niche communities, technological adoption, or speculative interest.
@@ V-D Spatial Clustering
Figure 4: Global clustering of countries based on decentralized finance-related tweets, sentiment scores, and economic indicators. Countries are color-coded according to their assigned cluster, which reflects similarities in their level of engagement with decentralized finance, public sentiment, and economic development. The classification is derived using a K-means clustering algorithm on 145 countries.
Figure 5: Scatter plot of GDP per capita, sentiment score, and proportion of tweets related to decentralized finance. Each point represents a country, positioned according to its GDP per capita and the proportion of relevant tweets, with color intensity reflecting sentiment score. The visualization highlights patterns in economic development, online engagement with decentralized finance, and public sentiment across different countries.
Figure 6: Box plot of GDP per capita, sentiment score, and proportion of tweets related to decentralized finance for each country cluster. The figure provides a statistical summary of the three key variables used in the clustering analysis, allowing for comparison between the three identified country clusters. The distribution of GDP per capita shows significant variation across clusters, with Cluster 3 containing the highest-income countries. The sentiment scores and tweet proportions also exhibit distinct trends, illustrating differences in engagement with decentralized finance across economic levels.
@@
Variable Cluster 1 Cluster 2 Cluster 3
GDP per Capita (USD)
Mean 3,075 4,222 33,879
Median 2,256 2,748 26,785
Min 240 399 4,099
Max 11,795 20,064 117,302
Std Dev 2,811 4,203 25,430
Tweet Proportion (%)
Mean 0.497 0.102 0.280
@@
Std Dev 0.470 0.075 0.186
Sentiment Score
Mean 0.706 0.651 0.665
@@ V-D1 Descriptive Statistics
Countries in Cluster 1 and Cluster 2 are characterized by relatively underdeveloped economies, with average GDP per capita values of \$3,075 and \$4,222, respectively. Our regression results in Table II indicate that countries with more advanced economic development tend to have higher public attention towards decentralized finance. However, countries in Cluster 1 present an interesting anomaly, exhibiting the highest level of attention to decentralized finance on social media, as reflected by the average proportion of related tweets (0.497%) and the most positive sentiment (average sentiment score of 0.706). In contrast, countries in Cluster 2 demonstrate the lowest level of attention to decentralized finance (average proportion of related tweets of 0.102%) and a relatively low sentiment (average sentiment score of 0.651).
Countries in Cluster 3, which have the most developed economies (average GDP per capita of \$33,879), show levels of attention to decentralized finance (average proportion of related tweets of 0.280%) and sentiment scores (average sentiment score of 0.665) that are lower than those in Cluster 1 but higher than those in Cluster 2. Figure 6 presents box plots depicting various countries based on the three variables mentioned above.
@@ V-D2 Crypto Legal Status
Considering the varying legal statuses of cryptocurrency trading in different countries, which may influence the level of attention and sentiment towards topics related to decentralized finance, we incorporated the ”Crypto Legal Status” data from the ”Cryptocurrency Regulation Tracker”. Box plots (Figure 7A) reveal a correlation between GDP per capita and the legality of cryptocurrencies. Excluding the outliers, South Korea and China, countries that recognize the legality of cryptocurrencies, those with a partial ban, and those with a general ban on cryptocurrencies show decreasing levels of economic development, with average GDP per capita values of \$24,009, \$7,635, and \$3,565, respectively. We assigned values to the three crypto legal statuses (Legal, Partial Ban, General Ban) as 0, 0.5, and 1, respectively, to represent the degree of prohibition of cryptocurrencies. The Spearman correlation coefficient between this variable and GDP per capita is -0.530 (p-value $<$ 0.001). The proportions of countries within different clusters that recognize the legality of cryptocurrency (Figure 7B) also show a similar pattern. In the most economically developed Cluster 3, 30% (15 out of 50) of the countries recognize the legality of cryptocurrency trading. In contrast, only 15% (5 out of 33) in Cluster 1 and 11% (7 out of 62) in Cluster 2 do so.
Figure 7: Cryptocurrency legal status across clusters. (A) Box plots illustrate the relationship between GDP per capita and the legal status of cryptocurrency trading in different countries. Countries that fully recognize cryptocurrencies tend to have the highest GDP per capita, while those with a partial or general ban exhibit decreasing economic development. (B) The bar chart presents the proportion of countries in each cluster that recognize the legality of cryptocurrencies. The most economically developed cluster (Cluster 3) has the highest percentage of countries where cryptocurrency trading is legal, whereas Clusters 1 and 2, with lower GDP per capita, exhibit greater regulatory restrictions.
Figure 8: Temporal evolution of topic probabilities for each cluster. This figure illustrates how discussions around decentralized finance have evolved across different economic clusters over the years. The topic probability trends reveal distinct thematic preferences among countries with different economic backgrounds. Countries in Cluster 1, despite their lower GDP per capita, exhibit sustained interest in specific aspects of decentralized finance, while mid- and high-GDP clusters display more varied topic distributions over time.
Figure 9: Comparison of shared and unique topics in decentralized finance discussions across clusters. This figure presents a Venn diagram-style analysis of the most frequently discussed topics within each cluster. While all clusters share a core interest in decentralized finance, their discussions diverge significantly in focus. Lowest-GDP countries (Cluster 1) emphasize financial hope and speculative gains, mid-GDP countries (Cluster 2) associate decentralized finance with social and religious aspects, and highest-GDP countries (Cluster 3) discuss cryptocurrencies primarily as investment and entertainment tools. Despite these distinctions, common themes such as financial innovation and blockchain adoption remain prevalent across all groups.
@@
Both lowest-GDP and mid-GDP countries emphasize the use of cryptocurrency to improve their economic situations, but the lowest-GDP countries surprisingly may consider cryptocurrency as a tool to achieve sudden wealth, characterized by words like ”luck”, ”hope”, and ”promote”. The mid-GDP countries focus on connecting social and religious elements with decentralized finance, using terms such as ”people”, ”god”, and ”lord”. However, the highest-GDP countries regard cryptocurrencies as novel financial tools to entertain and supplement daily life.